The long-distance calls made by South Africans are normally distributed with a mean of 16.3 minutes and a standard deviation of 4.2 minutes. What percentage of calls last more than 18 minutes (round off to two decimal places)?


1
Expert's answer
2020-10-05T17:42:33-0400

Let "X" be a random variable, which denotes the call duration in minutes: "X\sim N(\mu, \sigma ^2)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)."

Given "\mu=16.3\ min, \sigma=4.2\ min"


"P(X>18)=1-P(X\leq18)="

"=1-P(Z\leq\dfrac{18-16.3}{4.2})\approx1-P(Z\leq0.4047619)\approx"

"\approx1-0.657174=0.342826"

"34.28\%"